A biochemist is investigating the conversion of atmospheric nitrogen to usable form by bacteria that inhabit the root systems of certain legumes and needs to know the pressure in kiloPascals exerted by 1.25g of nitrogen gas in a flask of volume 500 mL at 37 0C.


1
Expert's answer
2021-02-25T03:20:17-0500

pV=nRT

p = nRT/V

n = m/M

M (N2) = 28 g/mol

R = 8.314 L kPa K−1 mol−1

n (N2) = 1.25/28 = 0.04 mol

p (N2) = (0.04 x 8.314 x (37 + 273))/ 0.5 = 206 kPa